OpenStudy (anonymous):

PLEASE HELP i'LL GIVE A MEDAL!!! What is the slope of the line that passes through the points (3, –1) and (–2, –5)? CHOICES LISTED BELOW

OpenStudy (anonymous):

OPTION 1: \[\frac{ 4 }{ 5 }\] OPTION 2: \[\frac{ 5 }{ 4 }\] OPTION 3: \[-\frac{ 4 }{ 5 }\] OPTION 4: \[-\frac{ 5 }{ 4 }\]

OpenStudy (anonymous):

@andrewhaze @AriPotta

OpenStudy (dtan5457):

\[\frac{ y2-y1 }{ x2-x1 }\] 3,-1 -2,-5 plug the points in, what is your slope?

OpenStudy (anonymous):

is it \[\frac{ 5 }{ 4 }\]

OpenStudy (andrewhaze):

well im really good at graphs so let me look

OpenStudy (dtan5457):

no it is not

OpenStudy (dtan5457):

3,-1 -2,-5 y2-y1=-5-(-1)=-4 x2-x1=-2-3=-5 -4/-5=4/5

OpenStudy (anonymous):

Thank you @dtan5457

OpenStudy (andrewhaze):

A